OpenStudy (anonymous):

Find the area that lies between the two circles r = 10sin(x) r = 10cos(x) This is a calc 2 polar question

OpenStudy (anonymous):

This is what it looks like 

OpenStudy (anonymous):

I was going to use this equation \[10\int\limits_{0}^{2\pi}\sin(\theta)d \theta+10\int\limits_{\pi/4}^{\pi/2}\cos(\theta)d \theta\]

OpenStudy (anonymous):

I know for polar you use this equation for area:\[\int\limits_{a}^{b} \pi (f(r))^2 d \theta\]

OpenStudy (anonymous):

I was not sure if I should use that here